OpenStudy (anonymous):

James has decided to start saving for a cruise over the summer. His money is currently in the local bank, modeled by the function s(x) = 102. He is able to do work around the neighborhood to earn extra money and that function would be a(x) = 8(x + 2), where x is measured in hours. Explain to James how he can create a function that combines the two and describe any simplification that can be done.

OpenStudy (hoblos):

a(x) = 8(x+2) = 8x + 16 add the two functions f(x) = s(x) + a(x) = 102 + 8x + 16 = 8x + 118

OpenStudy (anonymous):

Thank you so much, would you mind helping me with another?

OpenStudy (hoblos):

sure i'll try to help

OpenStudy (anonymous):

Victoria has some money from her birthday and the amount is modeled by the function h(x) = 200. She read about a bank that has savings accounts that accrue interest according to the function s(x) = (1.05)x − 1. Victoria is thinking about putting her money into the savings account to gain interest. Using complete sentences, explain to Victoria how she can combine her functions to create a new function, and explain what this new function means.

OpenStudy (hoblos):

well i'm not sure about it :S paste it as a new question so that if anybody else knows will try to help you